OpenStudy (anonymous):

Will FAN AND MEDAL!!! 2569+1+2+1+2+1+2+1+2+1+2+1+2+1+2+1+2+1+2+1+2+1+2+1+2*0=

OpenStudy (anonymous):

0

OpenStudy (anonymous):

0

OpenStudy (anonymous):

0

OpenStudy (mathstudent55):

2603

OpenStudy (anonymous):

2603, using google. Note: multiplication comes first—then addition.

OpenStudy (mathstudent55):

The zero is only multiplying the last 2. 2 * 0 = 0. All the other numbers must be added to 2569. 2569+1+2+1+2+1+2+1+2+1+2+1+2+1+2+1+2+1+2+1+2+1+2+1+2*0 = 2569+1+2+1+2+1+2+1+2+1+2+1+2+1+2+1+2+1+2+1+2+1+2+1+ 0 = 2569+1+2+1+2+1+2+1+2+1+2+1+2+1+2+1+2+1+2+1+2+1+2+1 = 2569+(1+2+1+2+1+2+1+2+1+2+1+2+1+2+1+2+1+2+1+2+1+2+1) = 2569 + 34 = 2603

OpenStudy (anonymous):

@mathstudent55 Thank you

OpenStudy (mathstudent55):

This would be zero: (2569+1+2+1+2+1+2+1+2+1+2+1+2+1+2+1+2+1+2+1+2+1+2+1+2)*0= In this case, you'd have to add all the numbers in the parentheses first, then multiply the sum by 0. The answer is zero. (2569+1+2+1+2+1+2+1+2+1+2+1+2+1+2+1+2+1+2+1+2+1+2+1+2) * 0= = 2605 * 0 = 0

OpenStudy (mathstudent55):

You're welcome.

OpenStudy (anonymous):

Thank you.